QUESTION THREE

a. The Y-axis intercept of the Security Market Line (SML) indicates the required rate of return on an individual stock with a beta of 1.0.

i. True

ii. False

b. State which type of risk each of the following are:

i. Labour strikes

ii. Rising inflation

iii. Equipment failure

iv. Management incompetence Currency BID/ Buying Ask/ Selling US Dollar $ 10 11 GB Pound £ 14 15 Euro € 11 12 Rand R 0.6 0.7 4

c. Using the table below, calculate the following for both stocks:

State of Economy Probability Returns of Stock A Returns of Stock B
Fair 0.4 15% 12%
Great 0.3 25% 19%
Poor 0.3 -10% 8%

i. What are the expected returns and standard deviations for these stocks?

ii. Using the information in the previous problem, suppose you have K 20, 000 in total. If you put K 15, 000 in Stock A and the remainder in Stock B,

what would be the expected return, covariance and standard deviation of your portfolio?

d. Explain the function of the SML.   

e. If the risk-free rate is 8 percent, the expected return on the market is 13 percent, and the expected return on Security J is 15 percent, then what is the beta of Security J?
Answer #1

a.FALSE (The Y-axis intercept of the Security Market Line (SML) indicates the required rate of return on an individual stock with a beta of 0 which is the risk free rate of return.)

b.i. Labour Strikes : It is a type of unsystematic risk. Unsystematic risks are unique to a particular firm or industry.

ii.Rising Inflation : It is a type of systematic risk. Such risks are related to the entire market and are unontrollable.

iii.Equipment failure: Based on the above explanation this is an unsystematic risk.

iv. Mnagement Incompetence : Based on the above explanation this is an unsystematic risk.

c.i. EXPECTED RETURN :

STOCK A = (0.4*15) + (0.3*25) + (0.3* (-10)) = 10.5%

STOCK B = (0.4*12) + (0.3*19) + (0.3* 8) = 9.9%

STANDARD DEVIATION :

STOCK A = square root of [(15-10.5)*0.4 + (25-10.5)*0.3 + (-10-10.5)*0.3] = 0

STOCK B = square root of [(12-9.9)*0.4 + (19-9.9)*0.3 + (8-9.9)*0.3] = square root of 30 = 5.477   

ii. Expected Return : = (15000*10.5%) + (5000*9.9%) = K 2070

Covariance :

Probability(1) Deviation A(2) Deviation B(3) (2)*(3)=(4) (1)*(4)

0.4 15-10.5 12-9.9 9.45 3.78

0.3 25-10.5 19-9.9 131.95 39.585

0.3 -10-10.5 8-9.9 38.95 11.685

covariance = 3.78 + 39.585 + 11.685 = 55.05

d. Function of SML :

Security Market Line reflects the relation between CAPM return and Beta.

It is the graphical representation of CAPM graph.